# Lie symmetry analysis of the two-Higgs-doublet model field equations

### v1

@@ 2.3 Symmetries of the action $\mathcal{S}$
since they (if Lie-type) generate Noether currents, and usually lifts to the quantized theory. As mentioned in Section 1, a transformation is called a variational symmetry of a theory, if the transformation leaves the action integral of the theory invariant,
@@
where $\beta^{\mu}$ is a local [^4] The $\beta$’s will be functions on a jet space with formal coordinates $(x,y,\ldots,y^{(n)})\in\mathbb{R}^{m}$ for some $n$ and $m$, hence non-local expressions like $\int ydx$ cannot occur in the $\beta$’s [35]. function of $x$, $y$ and derivatives of $y$ up to some order, and where $d_{\mu}\equiv{d}/{dx^{\mu}}$ is a total derivative. If we can choose $\beta^{\mu}=0$ for all $\mu$, and hence
@@
for a small parameter $\epsilon$, where $dF_{\mu}$ are the components of the differential outward normal vector of the boundary $\partial\Omega$ of $\Omega$. Then $\mathcal{S}$ is invariant if and only if the boundary term vanishes,
@@
which will be the case if the fields vanish sufficiently fast at infinity or are obeying certain periodic properties. Note that this boundary term does not depend on variations of the fields, so it is not automatically set to zero, unlike the variations at the boundary. Furthermore, it could be tempting to apply the criterion (2.20) directly to find all divergence symmetries, but it typically will lead us to very difficult, if not unsolvable, differential equations, and we are advised to proceed via the Euler-Lagrange equations, see e.g. [46].
A Lagrangian $\mathcal{L}^{\prime}=\mathcal{L}+\epsilon d_{\mu}\beta^{\mu}$ will yield the same Euler-Lagrange equations as $\mathcal{L}$, since a total divergence (also denoted a null Lagrangian) always will produce trivial Euler-Lagrange equations. In this article, the system of PDEs $\Delta=0$ will be a set of Euler-Lagrange equations of a Lagrangian $\mathcal{L}$. We define the Euler operator $E=(E_{1},\ldots,E_{q})$ with components given by formal, infinite series
@@
with $|J|$, $y^{i}_{J}$ and $D_{J}$ given by (2.9), (2.10) and (2.12), respectively. Then, the Euler-Lagrange equations of a Lagrangian $\mathcal{L}$ may be written
@@
A direct calculation show that each $E_{i}$ annihilate any total divergence $d_{\mu}\beta^{\mu}$, as implied above. In fact, it can be shown that a function $f$ is a total divergence if and only if it is annihilated by the Euler operator [35],
@@
The strict variational Lie point symmetries, given by the symmetry algebra $\mathfrak{g}_{\text{svar}}$, are obviously variational. The general variational Lie point symmetries, including divergence symmetries, will also generate a Lie symmetry algebra, which we denote $\mathfrak{g}_{\text{var}}$. Finally, all Lie point symmetries of the Euler-Lagrange equations of $\mathcal{L}$ will generate a Lie symmetry algebra $\mathfrak{g}_{\text{EL}}$, which will contain the two other algebras, since it can be shown that all variational symmetries are symmetries of the Euler-Lagrange equations of the theory (the converse does not hold in general). Hence, the following symmetry algebra inclusions always hold,
@@
Symmetries of $\mathfrak{g}_{\text{EL}}$ that are not included in the subalgebra $\mathfrak{g}_{\text{var}}$ are the non-variational Lie point symmetries. In section 4 we will demonstrate that for the 2HDM, the inclusions in (2.28) actually are equalities.
Due to Noether’s theorem, variational Lie point symmetries generate conserved currents. In case of a first-order Lagrangian $\mathcal{L}({x},{y},{y}^{(1)})$ a symmetry given by (2.6) induces a conservation law $d_{\mu}j^{\mu}=0$ with a conserved current
@@
Variational symmetries leaves the action invariant, and if the measure $\mathcal{D}y\equiv\prod_{i,x}dy^{i}(x)$ of $Z$ also is invariant under the transformation, the symmetry will be a symmetry of the quantized theory as well.
@@ 2.4 Theories with potentials
One of the goals of this work is to demonstrate how Lie symmetry analysis of PDEs can be applied to the Euler-Lagrange equations of particle physics models with potentials, such as multi-Higgs models. In this section we will show three results which may simplify the symmetry analysis of such models.
We start by defining the Fréchet derivative [35] of an $r$-tuple $P[y]\equiv P(x,y,\ldots,y^{(n)})$ of differential functions to be the operator $D_{P}$ which maps any $q$-tuple of differential (smooth) functions $U$ to
@@
Then the Fréchet derivative of a $r$-tuple of functions $P=(P_{1},\ldots,P_{r})$ is the $q\times r$ matrix with elements [35]
@@
The adjoint $D_{P}^{\ast}$ of the Fréchet derivative then is an operator with a similar property [35]
@@
where e.g. $D_{x}\cdot u(x)=u_{x}+uD_{x}$ on the right side of equation (2.33). Now, let $X$ be an infinitesimal generator with coefficients
@@
where the $\eta^{i}$’s are polynomials in the fields (i.e. the dependent variables). Then $X$ has characteristic
@@
where $\mathcal{J}$ is the Jacobian matrix of $\eta$, since only the $|J|=0$ term contributes in (2.33). Derivatives like $d_{\mu}$ and $\partial_{y^{i}_{,\mu}}$ do not commute. However, the following, useful commutation formula holds [35]
@@
be the Lagrangian of a theory where $T$ denotes the kinetic part (the sum of the kinetic terms), while $V(\varphi_{1},\ldots,\varphi_{m})$ is a potential, that is, a real polynomial where the variables $\varphi_{j}$ form a subset of the dependent variables of the theory,
$\varphi=\{\varphi_{1},\ldots,\varphi_{m}\}\subset y=\{y^{1},\ldots,y^{q}\},$ (2.40)
where we for simplicity will (re-)order the $\varphi$’s first in $y$,
@@
Note that $\varphi$ may consist of any of the fields of $y$, although the intention is to let $V(\varphi)=V(\phi)$, i.e. let $V$ be a scalar potential. Also note that the ”kinetic” part $T$ here actually may consist of any terms complementary to $V$. Moreover, we assume $E(L)=0$ does not imply any polynomial consequences of the form
@@
where $p$ is a non-trivial, multivariable polynomial in the same fields as the potential VV. [^5] In differential algebra language, $I(E(\mathcal{L}))\cap\mathbb{R}[\varphi_{1},\ldots,\varphi_{m}]=\{0\}$, where $I(E(\mathcal{L}))$ is the differential ideal generated by the Euler-Lagrange expressions $E(\mathcal{L})$ (it encodes all formal consequences of the Euler-Lagrange-equations), see e.g. [47], and $\mathbb{R}[\varphi_{1},\ldots,\varphi_{m}]$ is the ring of all polynomials in the variables $\varphi$ with real coefficients. This means we cannot derive any relation of the form (2.42), from the Euler-Lagrange equations. We will call such a theory a polynomial potential theory. Practically any multi-Higgs model will be an example of a polynomial potential theory, including the 2HDM. One reason is that each of the Euler-Lagrange equations $E_{i}(\mathcal{L})=0$ for multi-Higgs models will include distinct, non-mixed order two derivatives of the field $y^{i}$ that do not occur in any other $E_{j}(\mathcal{L})=0$ for $j\neq i$, and hence cannot be eliminated to produce some polynomial consequence (2.42). We here assume the multivariable polynomial $V$ may include any terms of any degree in the fields $\varphi$, but in some cases we will regard potentials without linear terms (but a constant in $V$ is allowed). However, terms linear in the fields will in these cases not be forbidden among the kinetic terms. We will now prove the following, useful result about evolutionary symmetries of polynomial potential theories:
@@
The linearized symmetry condition (2.17) applied on the Euler-Lagrange equations $E(\mathcal{L})=0$ gives
@@
since $\operatorname{pr}X=\operatorname{pr}X_{Q}$, because the former already is in evolutionary form. Assume $\operatorname{pr}X(T)=d_{\mu}\beta^{\mu}$ for a possibly vanishing current $\beta$. Substituting (2.37) in (2.44) then yields
@@
since $E(d_{\mu}\beta^{\mu})=0$ cf. (2.26), and where $\mathcal{J}$ was the Jacobian matrix of $\eta$. Then, if $E(\mathcal{L})=0$, the first and last terms in (2.45) vanish, cf. (2.43), i.e. for any $i$
@@
since there are no derivatives in the potential. The polynomial equation (2.46) will also hold when $E(\mathcal{L})\neq 0$, since $E(\mathcal{L})=0$ does not imply any polynomial relations (i.e. consequences) between the fields of $V$, as $\mathcal{L}$ was a polynomial potential theory. But then $\operatorname{pr}X(V)=C$ for a constant $C$, and $C=0$ since all terms of
@@
We will apply Theorem 1 in section 4 to show that symmetries have to be strictly variational, without having to consider all the numerous, different conditions on the parameters of the potential.
@@ 2.5 Scalar, variational symmetries in any spacetime dimension
Covariant derivatives, the gauge boson field strength tensors and the Higgs doublets are given by
@@
$\Phi_{j}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\phi_{4(j-1)+1}+i\phi_{4(j-1)+2}\\ \phi_{4(j-1)+3}+i\phi_{4(j-1)+4}\end{pmatrix}$ (2.55)
where $\sigma^{a}$ are the Pauli matrices, $1\leq a\leq 3$. We here also include the case where $K=0$, i.e. where the theory is a pure NHDM with no gauge singlets.
Then, the variational, scalar symmetries of theories given by $\mathcal{L}_{d}$ are independent of $d$: {restatable}propositionMainProp The variational Lie point symmetries transforming only scalars are the same for all NHDM+KS Lagrangians $\mathcal{L}_{d}$, regardless of the spacetime dimension $d\in\mathbb{N}$.
@@
A proof is given in Appendix A. ∎
@@
To find all scalar, variational Lie point symmetries of an NHDM+KS Lagrangian $\mathcal{L}_{4}$, it is enough to consider the simplified Lagrangian $\mathcal{L}_{1}$, since such symmetries are exactly the same for the two theories.
The simplified Lagrangian $\mathcal{L}_{1}$ will only contain 4 gauge fields and one free variable, and is hence computationally much less demanding than $\mathcal{L}_{4}$ containing 16 gauge fields and 4 free variables. The validity of Proposition 2.5 is confirmed by trying out the analysis of Section 4 on the $d=2$ variant of the 2HDM Lagrangian, $\mathcal{L}_{2}$, in addition to the $d=4$ 2HDM Lagrangian $\mathcal{L}_{4}$. [^6] The even simpler case $d=1$ generates some SYM-technical problems, since the kinetic terms of the gauge bosons vanishes in this case. However, we could have added arbitrary dummy-derivatives to circumvent this, since the kinetic gauge terms are annihilated by scalar symmetries $\operatorname{pr}X$, and hence become irrelevant for the determination of scalar symmetries. After specializing to scalar transformations, the two approaches yield exactly the same equations, and hence exactly the same symmetry algebras (all Lie symmetries of the 2HDM turn out to be strictly variational and will hence be shared by the two Lagrangians according to Proposition 2.5). However, in the $d=2$ case the number of free variables and the number of gauge fields are halved, i.e. the field equations
@@
form a system of 16 equations of 16 dependent and 2 free variables, while the field equations
@@
form a system of 24 equations of 24 dependent and 4 free variables. The computation time for SYM to calculate the determining equations was reduced to a tiny fraction in the $d=2$ case. [^7] For the 2HDM the calculation of the determining equations of $\mathcal{L}_{4}$ took nearly 14 hours on a desktop computer, while the corresponding calculation for $\mathcal{L}_{2}$ on the same computer took only 45 minutes.

### v2

@@ 2.3 Symmetries of the action $\mathcal{S}$
since they, if they are of Lie type, generate Noether currents and are usually preserved in the quantized theory. As mentioned in Section 1, a transformation is called a variational symmetry of a theory, if the transformation leaves the action integral of the theory invariant,
@@
where $\beta^{\mu}$ is a local [^4] The $\beta$’s will be functions on the jet space $J^{n}$, hence non-local expressions such as $\int ydx$ cannot occur in the $\beta$’s [35]. function of $x$, $y$ and derivatives of $y$ up to some order, and where $d_{\mu}\equiv{d}/{dx^{\mu}}$ is a total derivative. If we can choose $\beta^{\mu}=0$ for all $\mu$, and hence
@@
for a small parameter $\epsilon$, where $dF_{\mu}$ are the components of the differential outward normal vector of the boundary $\partial\Omega$ of $\Omega$. Then, $\mathcal{S}$ is invariant if and only if the boundary term vanishes,
@@
which will be the case if the fields vanish sufficiently fast at infinity or obey certain periodic properties. Note that this boundary term does not depend on variations of the fields; therefore, it is not automatically zero, unlike variations at the boundary. Furthermore, it could be tempting to apply the criterion (2.22) directly to find all divergence symmetries, but it typically will lead us to very difficult, if not unsolvable, differential equations, and we are advised to proceed via the Euler-Lagrange equations [44].
A Lagrangian $\mathcal{L}^{\prime}=\mathcal{L}+\epsilon d_{\mu}\beta^{\mu}$ will yield the same Euler-Lagrange equations as $\mathcal{L}$, since a total divergence (also denoted a null Lagrangian) will always produce trivial Euler-Lagrange equations. In this study, the system of PDEs $\Delta=0$ is a set of Euler-Lagrange equations of a Lagrangian $\mathcal{L}$. We define the Euler operator $E=(E_{1},\ldots,E_{q})$ with components given by a formal infinite series
@@
where $|J|$, $y^{i}_{J}$ and $D_{J}$ are given by (2.11), (2.12) and (2.14), respectively. Then, the Euler-Lagrange equations of a Lagrangian $\mathcal{L}$ can be written as
@@
A direct calculation shows that each $E_{i}$ annihilates any total divergence $d_{\mu}\beta^{\mu}$, as previously implied. In fact, it can be shown that a function $f$ is a total divergence if and only if it is annihilated by the Euler operator [35],
@@
The strict variational Lie point symmetries, given by the symmetry algebra $\mathfrak{g}_{\text{svar}}$, are obviously variational. The general variational Lie point symmetries, including divergence symmetries, also generate a Lie symmetry algebra, which we denote $\mathfrak{g}_{\text{var}}$. Finally, all Lie point symmetries of the Euler-Lagrange equations of $\mathcal{L}$ will generate a Lie symmetry algebra $\mathfrak{g}_{\text{EL}}$, which will contain the two other algebras, because it can be shown that all variational symmetries are symmetries of the Euler-Lagrange equations of the theory (the converse does not hold in general). Hence, the following symmetry algebra inclusions always hold:
@@
The symmetries of $\mathfrak{g}_{\text{EL}}$ that are not included in the subalgebra $\mathfrak{g}_{\text{var}}$ are non-variational Lie point symmetries. In section 4 we will demonstrate that for the 2HDM, the inclusions in (2.30) actually are equalities.
According to Noether’s theorem, variational Lie point symmetries generate conserved currents. In the case of a first-order Lagrangian $\mathcal{L}({x},{y},{y}^{(1)})$ a symmetry given by (2.8) induces a conservation law $d_{\mu}j^{\mu}=0$ with a conserved current
@@
Variational symmetries leave the action invariant, and if the measure $\mathcal{D}y\equiv\prod_{i,x}dy^{i}(x)$ of $Z$ is also invariant under the transformation, the symmetry will also be a symmetry of the quantized theory.
@@ 2.4 Theories with potentials
One of the goals of this work is to demonstrate how Lie symmetry analysis of PDEs can be applied to the Euler-Lagrange equations of particle physics models with potentials, such as multi-Higgs models. In this section we present three results which may simplify the symmetry analysis of such models.
We start by defining the Fréchet derivative [35] of an $r$-tuple $P[y]\equiv P(x,y,\ldots,y^{(n)})$ of differential functions as the operator $D_{P}$ that maps any $q$-tuple of differential (smooth) functions $U$ to
@@
Then, the Fréchet derivative of an $r$-tuple of functions $P=(P_{1},\ldots,P_{r})$ is the $r\times q$ matrix with elements [35]
@@
The adjoint $D_{P}^{\ast}$ of the Fréchet derivative is an operator with similar properties [35],
@@
where e.g. $D_{x}\cdot u(x)=u_{x}+uD_{x}$ on the right-hand side of equation (2.35). Let $X$ be an infinitesimal generator with coefficients
@@
where the $\eta^{i}$ are polynomials in the fields (i.e., the dependent variables). Then, $X$ has characteristic
@@
where $\mathcal{J}$ is the Jacobian matrix of $\eta$, since only the $|J|=0$ term contributes to (2.35). Derivatives such as $d_{\mu}$ and $\partial_{y^{i}_{,\mu}}$ do not commute. However, the following, useful commutation formula holds [35]
@@
be the Lagrangian density of a theory, where $T$ denotes the kinetic part (the sum of the kinetic terms), while $V(\varphi_{1},\ldots,\varphi_{m})$ is a potential, i.e. a real polynomial in a subset of the dependent variables. Specifically, let
$\varphi=\{\varphi_{1},\ldots,\varphi_{m}\}\subset\{y^{1},\ldots,y^{q}\}=y.$ (2.42)
For convenience, we assume that the dependent variables are ordered such that
@@
An $n$th-order Lagrangian $\mathcal{L}$ may be viewed as a function on the jet space $J^{n}$, $\mathcal{L}:J^{n}\to\mathbb{R}$. Note that $\varphi$ may consist of any of the fields of $y$, although the intention is to let $V(\varphi)=V(\phi)$, that is, let $V$ be a scalar potential. Also note that the ”kinetic” part $T$ here actually may consist of any terms complementary to $V$. Moreover, we assume $E(L)=0$ does not imply any polynomial consequences of the form
@@
where $p$ is a non-trivial, multivariable polynomial in the same fields as the potential VV. [^5] In differential algebra language, $I(E(\mathcal{L}))\cap\mathbb{R}[\varphi_{1},\ldots,\varphi_{m}]=\{0\}$, where $I(E(\mathcal{L}))$ is the differential ideal generated by the Euler-Lagrange expressions $E(\mathcal{L})$ (it encodes all formal consequences of the Euler-Lagrange-equations), see for example [45], and $\mathbb{R}[\varphi_{1},\ldots,\varphi_{m}]$ is the ring of all polynomials in the variables $\varphi$ with real coefficients. This means that we cannot derive any relation of the form (2.44), from the Euler-Lagrange equations. We will call such a theory a polynomial potential theory. Practically any multi-Higgs model is an example of a polynomial potential theory, including the 2HDM. One reason is that each Euler-Lagrange equation $E_{i}(\mathcal{L})=0$ in multi-Higgs models will include distinct, second-order derivatives of the field $y^{i}$ that do not appear in any other $E_{j}(\mathcal{L})=0$ for $j\neq i$, and hence cannot be eliminated to produce some polynomial consequence (2.44). Here, we assume the multivariable polynomial $V$ may include any terms of any degree in the fields $\varphi$, but in some cases we will regard potentials without linear terms (but a constant in $V$ is allowed). However, in these cases, the linear terms are not forbidden among the kinetic terms. We now prove the following, useful result regarding evolutionary symmetries of polynomial potential theories:
@@
The linearized symmetry condition (2.19) applied on the Euler-Lagrange equations $E(\mathcal{L})=0$ yields
@@
since $\operatorname{pr}X=\operatorname{pr}X_{Q}$, because the former is already in evolutionary form. Assume
$\operatorname{pr}X(T)=d_{\mu}\beta^{\mu},$ (2.47)
for a possible vanishing current $\beta$. Substituting (2.39) into (2.46) then yields,
@@
because $E(d_{\mu}\beta^{\mu})=0$ cf. (2.28), and where $\mathcal{J}$ is the Jacobian matrix of $\eta$. Then, if $E(\mathcal{L})=0$, the first and last terms in (2.48) vanish, cf. (2.45), that is, for any $i$
@@
because there are no derivatives in the potential. The polynomial equation (2.49) will also hold when $E(\mathcal{L})\neq 0$, because $E(\mathcal{L})=0$ does not imply any polynomial relations (i.e., consequences) between the fields of $V$, as $\mathcal{L}$ is a polynomial potential theory. But then $\operatorname{pr}X(V)=C$ for constant $C$, and $C=0$ because all terms of
@@
We will apply Theorem 1 in Section 4 to demonstrate that symmetries must be strictly variational, without having to consider all the numerous, different conditions on the parameters of the potential.
@@ 2.5 Scalar, variational symmetries in any spacetime dimension
The covariant derivatives, gauge boson field strength tensors and Higgs doublets are given by:
@@
$\Phi_{j}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\phi_{4(j-1)+1}+i\phi_{4(j-1)+2}\\ \phi_{4(j-1)+3}+i\phi_{4(j-1)+4}\end{pmatrix}$, (2.58)
where $\sigma^{a}$ are the Pauli matrices, $1\leq a\leq 3$ and $Y$ denotes the $\mathsf{U}(1)_{Y}$ hypercharge of the corresponding field. Geometrically, $W^{a}_{\mu}$ can be viewed as the local components of an $\mathsf{SU}(2)_{L}$ principal connection, while $B_{\mu}$ are the local components of a $\mathsf{U}(1)_{Y}$ principal connection. Here, we also include the case where $K=0$, that is, where the theory is a pure NHDM with no gauge singlets.
Then, the variational, scalar symmetries of theories given by $\mathcal{L}_{d}$ are independent of $d$: {restatable}propositionMainProp The variational Lie point symmetries transforming only the scalars are the same for all NHDM+KS Lagrangians $\mathcal{L}_{d}$, regardless of the spacetime dimension $d\in\mathbb{N}$.
@@
The proof of this is provided in Appendix A. ∎
@@
To find all scalar, variational Lie point symmetries of an NHDM+KS Lagrangian $\mathcal{L}_{4}$, it is sufficient to consider the simplified Lagrangian $\mathcal{L}_{1}$, because such symmetries are exactly the same for the two theories.
The simplified Lagrangian $\mathcal{L}_{1}$ will only contain four gauge fields and one free variable and is hence computationally much less demanding than $\mathcal{L}_{4}$ that contains 16 gauge fields and four free variables. The validity of Proposition 2.5 is confirmed by performing the analysis of Section 4 on the $d=2$ variant of the 2HDM Lagrangian $\mathcal{L}_{2}$, in addition to the $d=4$ 2HDM Lagrangian $\mathcal{L}_{4}$. [^6] The even simpler case $d=1$ causes some SYM-related technical issues, because the kinetic terms of the gauge bosons vanish in this case. However, we could have added arbitrary dummy derivatives to circumvent this, as the kinetic gauge terms are annihilated by scalar symmetries $\operatorname{pr}X$, and therefore become irrelevant for determining scalar symmetries. After specializing to scalar transformations, the two approaches yield exactly the same equations, and hence exactly the same symmetry algebras (all Lie symmetries of the 2HDM turn out to be strictly variational and will hence be shared by the two Lagrangians according to Proposition 2.5). However, in the $d=2$ case, the number of free variables and gauge fields are halved, that is, the field equations
@@
form a system of 16 equations with 16 dependent and two free variables, whereas the field equations
@@
form a system of 24 equations with 24 dependent and four free variables. The computation time for SYM to calculate the determining equations was reduced to a small fraction in the $d=2$ case. [^7] For the 2HDM the calculation of the determining equations of $\mathcal{L}_{4}$ took nearly 14 hours on a desktop computer, whereas the corresponding calculation for $\mathcal{L}_{2}$ on the same computer took only 45 minutes.